# Neural Automated Writing Evaluation with Corrective Feedback

### v1

@@ Abstract
Izia Xiaoxiao Wang Xihan Wu Edith Coates††thanks: Equally contributed authors. Min Zeng Jiexin Kuang Siliang Liu Mengyang Qiu Jungyeul Park Linguistik Zentrum Zürich, Universität Zürich, Schweiz Department of Linguistics, The University of British Columbia, Canada Department of Mathematics, The University of British Columbia, Canada Department of Psychology, Trent University, Canada
@@ 3 System Description
We build seq2seq GEC models using fairseq (Ott et al., 2019). We explore ensemble models that incorporate different pre-trained large language models (LLMs), including BERT and BART. The key difference between the two LLMs is in their pre-training objectives. BERT is trained using a masked language modelling objective, where a certain percentage of the input tokens are masked, and the model is trained to predict the original tokens. In contrast, BART is trained using a denoising auto-encoding objective, where the model is trained to reconstruct the original input text from a corrupted version of the input. Our current training setup is largely based on Kaneko et al. (2020). We use four GEC datasets – FCE, NUCLE, W&I +LOCNESS, and Lang-8 – provided by the BEA 2019 Shared Task for training (Bryant et al., 2019).444https://www.cl.cam.ac.uk/research/nl/bea2019st All the incorrect sentences are spell-checked and added back in order to augment the size of the original training set (1,157,324 sentence pairs), resulting in 1,718,693 sentence pairs in total. 6,575 sentence pairs from the FCE and W&I+LOCNESS development sets are used as validation. Our best model (using BERT) achieves a 65.29 F score on the BEA 2019 test set. We have also been exploring various ways of fine-tuning the power of LLMs as well as edit-based approaches such as GECToR (Omelianchuk et al., 2020) and T5 (Rothe et al., 2021), which currently have provided SOTA results for the GEC task. These pre-trained models are prepared for seamless integration into our system. While prompting GPT for the GEC task has also been been investigated, it could not obtain the state of the art results for the GEC task, and this is still prevalent in recent previous work (Loem et al., 2023).
@@ 4 Discussion and Future Perspectives
- On a scale of 1-5, how fair do you think the grammar feedback on this writing to be? (1= not fair at all, 5= very fair) |
- On a scale of 1-5, how accurate and trustworthy do you think the grammar correction on this writing to be? (1= not accurate/trustworthy at all, 5= very accurate/trustworthy) |
- On a scale of 1-5, how fair do you think the grading and evaluation of this writing to be? (1= not fair at all, 5= very fair) |
- On a scale of 1-5, how accurate and trustworthy do you think the grading and evaluation of this writing to be? (1= not accurate/trustworthy at all, 5= very accurate/trustworthy) |
@@
- On a scale of 1-5, how clear do you think the grading and feedback on this writing to be? (1= not clear at all, 5= very clear) |
- On a scale of 1-5, how helpful do you perceive the feedback on this writing to be in improving this student’s writing skills? (1= not helpful at all, 5= very helpful) |
- On a scale of 1-5, how helpful do you think rubrics with criteria and scores to be in understanding the grading and feedback on this writing assignment? (1= not helpful at all, 5= very helpful) |
@@
- On a scale of 1-5, how difficult do you find it to perceive the grading and feedback on this writing to be? |
- On a scale of 1-5, how much mental effort does this task (differ for stages 1 and 2) require from you to perceive the feedback? |
@@
- On a scale of 1-5, how comfortable do you feel about perceiving/using this system to receive the grading and feedback on this writing? |
- On a scale of 1-5, how effective do you feel about perceiving/using this system to receive the grading and feedback on this writing? |

### v2

@@ Abstract
Izia Xiaoxiao Wang1∗ Xihan Wu2∗ Edith Coates3††thanks: Equally contributed authors. Min Zeng2 Jiexin Kuang2 Siliang Liu2 Mengyang Qiu4 Jungyeul Park2 1Linguistik Zentrum Zürich, Universität Zürich, Schweiz 2Department of Linguistics, The University of British Columbia, Canada 3Department of Mathematics, The University of British Columbia, Canada 4Department of Psychology, Trent University, Canada
@@ 3 System Description
We build seq2seq GEC models using fairseq (Ott et al., 2019). We explore ensemble models that incorporate different pre-trained large language models (LLMs), including BERT and BART. The key difference between the two LLMs is in their pre-training objectives. BERT is trained using a masked language modelling objective, where a certain percentage of the input tokens are masked, and the model is trained to predict the original tokens. In contrast, BART is trained using a denoising auto-encoding objective, where the model is trained to reconstruct the original input text from a corrupted version of the input. Our current training setup is largely based on Kaneko et al. (2020). We use four GEC datasets – FCE, NUCLE, W&I +LOCNESS, and Lang-8 – provided by the BEA 2019 Shared Task for training (Bryant et al., 2019).444https://www.cl.cam.ac.uk/research/nl/bea2019st All the incorrect sentences are spell-checked and added back in order to augment the size of the original training set (1,157,324 sentence pairs), resulting in 1,718,693 sentence pairs in total. 6,575 sentence pairs from the FCE and W&I+LOCNESS development sets are used as validation. Our best model (using BERT) achieves a 65.29 F0.5 score on the BEA 2019 test set. We have also been exploring various ways of fine-tuning the power of LLMs as well as edit-based approaches such as GECToR (Omelianchuk et al., 2020) and T5 (Rothe et al., 2021), which currently have provided SOTA results for the GEC task. These pre-trained models are prepared for seamless integration into our system. While prompting GPT for the GEC task has also been been investigated, it could not obtain the state of the art results for the GEC task, and this is still prevalent in recent previous work (Loem et al., 2023).
@@ 4 Discussion and Future Perspectives
| - On a scale of 1-5, how fair do you think the grammar feedback on this writing to be? (1= not fair at all, 5= very fair) | |
| - On a scale of 1-5, how accurate and trustworthy do you think the grammar correction on this writing to be? (1= not accurate/trustworthy at all, 5= very accurate/trustworthy) | |
| - On a scale of 1-5, how fair do you think the grading and evaluation of this writing to be? (1= not fair at all, 5= very fair) | |
| - On a scale of 1-5, how accurate and trustworthy do you think the grading and evaluation of this writing to be? (1= not accurate/trustworthy at all, 5= very accurate/trustworthy) | |
@@
| - On a scale of 1-5, how clear do you think the grading and feedback on this writing to be? (1= not clear at all, 5= very clear) | |
| - On a scale of 1-5, how helpful do you perceive the feedback on this writing to be in improving this student’s writing skills? (1= not helpful at all, 5= very helpful) | |
| - On a scale of 1-5, how helpful do you think rubrics with criteria and scores to be in understanding the grading and feedback on this writing assignment? (1= not helpful at all, 5= very helpful) | |
@@
| - On a scale of 1-5, how difficult do you find it to perceive the grading and feedback on this writing to be? | |
| - On a scale of 1-5, how much mental effort does this task (differ for stages 1 and 2) require from you to perceive the feedback? | |
@@
| - On a scale of 1-5, how comfortable do you feel about perceiving/using this system to receive the grading and feedback on this writing? | |
| - On a scale of 1-5, how effective do you feel about perceiving/using this system to receive the grading and feedback on this writing? | |
@@ 5 Conclusion
Acknowledgement
This research is based upon work partially supported by the SoTL Seed Program for Jungyeul Park at The University of British Columbia.